The digits of a three digit number forms a Gp. If we subtract 400 From it, then we get another 3  Digit Number whose digits Forms an Arithmetic series. What is The Sum Of These two Numbers? 

Grade:12th pass

2 Answers

Arun
5 years ago
 
Dear student
 
Let the digits are x , y and z
 
hence number is 100z + 10 y + x
 
y^2 = xz Hence xz must be a square number
 
xz = 9
 
hence y = 3
 
now some hit and trial,
x = 1 , z= 9
 
hence number is 931
 
 
subtract 400,
931 –  400 = 531
 
hence 931 + 531 = 1462
Arun
5 years ago
 
Let the digits are x , y and z
 
hence number is 100z + 10 y + x
 
y^2 = xz Hence xz must be a square number
 
xz = 9
 
hence y = 3
 
now some hit and trial,
x = 1 , z= 9
 
hence number is 931
 
 
subtract 400,
931 –  400 = 531
 
hence 931 + 531 = 1462
 
Hope it helps
 
Regards
Arun (askIITians forum expert)